OpenStudy (hvploe_):

Karen has a part time job babysitting, where she earns $6 pe hour. She also walks dogs and earns $4 per dog that she walks. Her monthly earnings can be expressed with the equation 6x+4y=700, where x is the number of hours karen works and y is the number of dogs she walks

OpenStudy (opti):

Need a q

OpenStudy (hvploe_):

1)-You must change the equation that represents her chosen college student to slope intercept form. indentify the slope and y-intercept of the equation.be sure to show all your work.

OpenStudy (hvploe_):

You there?

OpenStudy (hvploe_):

Hello??

OpenStudy (opti):

solve for y

OpenStudy (hvploe_):

How?

OpenStudy (hvploe_):

Everything I get turns into a decimal

OpenStudy (opti):

show me

OpenStudy (hvploe_):

6x+4y=700 700/4=175/6=29.2

OpenStudy (hvploe_):

Hello?

OpenStudy (opti):

\[6x+4y=700\] \[4y=700 - 6x\] \[y=\frac{ 700 }{ 4 } - \frac{ 6x}{ 4 }\]

OpenStudy (hvploe_):

That still gives a decimal for 6y/4

OpenStudy (hvploe_):

175-1.5=173.5

OpenStudy (opti):

and

OpenStudy (hvploe_):

That's the slope intercept form?

OpenStudy (hvploe_):

I'm sorry I suck at this

OpenStudy (opti):

no one starts off good at math. its okay

OpenStudy (hvploe_):

So how can I transfer it into slope intercept form?

OpenStudy (guoalice8):

slope intercept form is y=ax+b

OpenStudy (opti):

I did it earlier